OpenStudy (anonymous):

Find the length of AC. AB=36cm

OpenStudy (anonymous):



OpenStudy (anonymous):

Well it is diameter and radius!

OpenStudy (anonymous):

I know diameter would be 36 and the radius would be 18

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

Any point to the middle of the circle is what?

OpenStudy (anonymous):

All the points on the circle to the center of the circle are the same length

OpenStudy (anonymous):

it would be the radius

OpenStudy (anonymous):

YUP! which is 18

OpenStudy (anonymous):

so then do i use the radius to find ac?

OpenStudy (anonymous):

No the radius is ac...

OpenStudy (anonymous):

I mean the diameter is ac

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

2pi*r*angle/360= lenght

OpenStudy (anonymous):

Because ap is 18 and pc is 18

OpenStudy (anonymous):

2*pi*18*150/360=AC

OpenStudy (anonymous):

This got confusing...Whats the name of that equation @cinar

OpenStudy (anonymous):

I always thought you used that to find arc lengths

OpenStudy (anonymous):

2*pi*r= circumstances of circle right? 2*pi*r*angle/360= arc length corresponding to the angle..

OpenStudy (anonymous):

2*pi*r*360/360= circumstances of circle =2*pi*r 2*pi*r*180/360=length of half circle=pi*r 2*pi*r*90/360=length of quarter=pi*r/2 2*pi*r*angle/360=length of arc corresponding to the angle..

OpenStudy (anonymous):

for example length of BC=2*pi*r*30/360=2*pi*18*30*360